# Quantifying Aleatoric and Epistemic Dynamics Uncertainty via Local Conformal Calibration

### v2

@@ 7.0.1 Acknowledgements
Appendix 0.A Proving Multi-Step Coverage with Linear Dynamics Approximations
For convenience, we define $\hat{\mathcal{N}}_{\tau+1}(\hat{\mu}_{\tau+1},\hat{\Sigma}_{\tau+1})=\tilde{f}_{\hat{\mu}_{\tau},u_{\tau}}(\hat{\Sigma}_{\tau})=\tilde{f}(\hat{\mathcal{N}}_{\tau}(\hat{\mu}_{\tau},\hat{\Sigma}_{\tau}),u_{\tau})$. Let $\xi_{\tau+1}\in\mathbb{R}$ denote the scaling factor required to calibrate the $\hat{\mathcal{N}}_{\tau+1}$ MVN, i.e. the minimum scalar such that the $(1-\alpha)$ confidence region of $\hat{\mathcal{N}}_{\tau+1}(\hat{\mu}_{\tau+1},\xi_{\tau+1}\hat{\Sigma}_{\tau+1})$ contains at least $(1-\alpha)\%$ of the true next states $s_{\tau+1}$. In principle, $\xi_{\tau+1}$ could be obtained by the procedure detailed in the main paper: 1) obtain a dataset $D_{cal}$ of exchangeable transition tuples $((s_{\tau},u_{\tau}),s_{\tau+1})$; 2) transform it into a dataset of exchangeable $((\hat{\mathcal{N}}_{\tau},u_{\tau}),s_{\tau+1})$ tuples through a fixed mapping, e.g. $g$ as defined in Equation (11); 3) propagate all state MVNs using $\tilde{f}_{\hat{\mu}_{\tau},u_{\tau}}$; 4) calculate the residuals, via Equation (12), between $\hat{\mathcal{N}}_{\tau+1}$ and $s_{\tau+1}$; 5) determine the $\hat{q}_{\tau+1}$ cutoff and $\xi_{\tau+1}$ using SplitCP (§$\S$4). By construction $\xi_{\tau+1}\geq 0$ and almost surely $\xi_{\tau+1}>0$, thus $\xi_{\tau+1}\hat{\Sigma}_{\tau+1}$ is still a valid covariance. For $\tau=t$, $\xi_{t+1}$ is the scaling factor defined in Theorem 5.2 for one-step prediction. There the fixed mapping $g(X^{(i)})=(\hat{\mathcal{N}}(s_{t}^{(i)},\Sigma_{0}),u_{t}^{(i)})$ made the MVNs constructed in step 2) have covariance $\hat{\Sigma}_{t}=\Sigma_{0}$. Unlike in the first planning step where the starting state uncertainty $\Sigma_{0}$ can be pre-defined by the user, for $\tau>t$ the pre-transition state uncertainty $\hat{\Sigma}_{\tau}=\hat{\Sigma}_{\tau-1,cal}$ can take a wide range of forms, depending on the preceding actions. For clarity, we will now use $\tau=\{t+1,\ldots,t+H-1\}$ to exclude the previous $\tau=t$ entry (as there $\xi_{\tau+1}=\xi_{t+1}$ directly). Since it is not tractable to redo steps 2-5) above for all possible pre-transition state covariances, we will show that if the approximate dynamics $\tilde{f}$ are linear then $\xi_{\tau+1}\leq\xi_{t+1}$. While there is no guarantee that $\hat{\Sigma}_{\tau}$ is exchangeable with $\bar{D}_{cal}$, and thus $\xi_{t+1}$ might not be the minimum required scaling factor, for linear $\tilde{f}$ using $\xi_{t+1}$ ensures the $(1-\alpha)$ coverage guarantee holds for the entire plan.
{assumption}
The approximate dynamics are linear, i.e. $\tilde{f}_{\hat{\mu}_{\tau},u_{\tau}}(\hat{\Sigma}_{\tau})=\hat{\mathcal{N}}_{\tau+1}(A\hat{\mu}_{\tau}+Bu_{\tau},A\hat{\Sigma}_{\tau}A^{\top}+Q)$ where $A,B,Q$ are matrices of appropriate dimension.
With linear $\tilde{f}$, $\hat{\mu}_{\tau+1}$ depends only on $\hat{\mu}_{\tau}$ and $u_{\tau}$, i.e. the mean of the approximate predictive distribution is independent of the initial state uncertainty.
Corollary 1.
For linear $\tilde{f}$, $\hat{\mu}_{\tau+1}=\mathbb{E}[\tilde{f}(\hat{\mathcal{N}}(\hat{\mu}_{\tau},\hat{\Sigma}_{\tau}),u_{\tau})]$ is independent of $\hat{\Sigma}_{\tau}$.
Hence, for fixed mappings from $D_{cal}$ to $\bar{D}_{cal}$ that set the same $\hat{\mu}_{\tau}$, the resulting $\hat{\mu}_{\tau+1}$ will also be the same irrespective of $\hat{\Sigma}_{\tau}$. We further require $\Sigma_{0}$ to be the "smallest" state uncertainty along the plan, which can be formalized by the following Loewner order [^2] The Loewner order is a partial ordering between hermitian matrices $M,N$ in which $M>N$ ($M\geq N$) means that $M-N>0$ ($M-N\geq 0$), i.e. their difference is a positive (semi-)definite matrix. See [19] for more information..
{assumption}
We assume that $\Sigma_{0}\leq\hat{\Sigma}_{\tau}$ for $\tau\in\{t+1,\ldots,t+H-1\}$. If $\Sigma_{0}$ is diagonal with entries of equal value (we use $\Sigma_{0}=2\cdot 10^{-5}I$), then its non-zero entries can be interpreted as a minimum bound on the eigenvalues of $\hat{\Sigma}_{\tau}$ so that Assumption 0.A holds. Practically, this implies that the $(1-\alpha)$ confidence region of $\Sigma_{0}$ is always a subset of the $(1-\alpha)$ confidence region of $\hat{\Sigma}_{\tau}$.
Corollary 2.
Let $\varepsilon_{\mathcal{N},\alpha}$ be the $(1-\alpha)$ confidence region of MVN $\mathcal{N}$, then ${\Sigma_{0}\leq\hat{\Sigma}_{\tau}}\Rightarrow\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},{\Sigma}_{0}),\alpha}\subseteq\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},\hat{\Sigma}_{\tau}),\alpha}$, $\forall\alpha\in(0,1)$.
@@ Appendix 0.A Proving Multi-Step Coverage with Linear Dynamics Approximations
Since $\Sigma_{0}\leq\hat{\Sigma}_{\tau}\Rightarrow\Sigma_{0}^{-1}\geq\hat{\Sigma}_{\tau}^{-1}$, then $z^{\top}\Sigma_{0}^{-1}z\geq z^{\top}\hat{\Sigma}_{\tau}^{-1}z$ for any vector $z$ of appropriate dimensions. Recalling that $(y-\mu)^{\top}\Sigma^{-1}(y-\mu)$ is the squared Mahalanobis distance, then by Lemma 2 $\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},\hat{\Sigma}_{0}),\alpha}\subseteq\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},\hat{\Sigma}_{\tau}),\alpha}$ since $y\in\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},\hat{\Sigma}_{0}),\alpha}\Rightarrow y\in\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},\hat{\Sigma}_{\tau}),\alpha}$. ∎
This is illustrated in Figure 5, along with a representation of steps 2 to 4. If $\tilde{f}_{\hat{\mu}_{\tau},u_{\tau}}$ preserves the Loewner order relationship after propagation, i.e. $\Sigma_{0}\leq\hat{\Sigma}_{\tau}\Rightarrow\hat{\Sigma_{1}}\leq\hat{\Sigma}_{\tau+1}$, we call it matrix monotone. Below we show that linear dynamics are matrix monotone, for any system ($A$) and input ($B$) matrices, if the noise matrix ($Q$) is positive semi-definite.
Figure 5: Visualizing Loewner order and its implication for multi-step planning. If we have the Loewner ordering $\hat{\Sigma}_{\tau}\geq\Sigma_{0}$ (following Assumption 0.A), then the $(1-\alpha)$ confidence ellipse of $\Sigma_{0}$ is a subset of the $(1-\alpha)$ confidence ellipse of $\hat{\Sigma}_{\tau}$, $\forall\alpha\in[0,1]$. The state-action pair $X^{(i)}$ (red) of $D_{cal}$ is converted into a state MVN and action tuple by a fixed mapping (step 2), either $g$ (green) or $h$ (pink). This is then propagated through $\tilde{f}$ (yellow) to obtain an estimate of the future robot state $\hat{\mathcal{N}}_{\tau+1}^{(i)}$ (step 3). If $\tilde{f}$ is linear, the Loewner ordering is preserved (see Lemma 3). From Corollary 1, the estimated state means (black) are the same for both mappings. Finally, the residuals $R^{(i)}$ (blue) are calculated between the estimated state MVNs and the true next state $s_{\tau+1}$ (step 4) show in red. We show in Theorem 0.A.1 that $R^{(i)}_{\tau+1}\leq R^{(i)}_{1},\forall i\in\{1,\ldots,\lvert D_{cal}\rvert\}$.
Lemma 3.
The linear transformation $\tilde{f}_{\hat{\mu}_{\tau},u_{\tau}}(\hat{\Sigma}_{\tau})=\hat{\mathcal{N}}_{\tau+1}(A\hat{\mu}_{\tau}+Bu_{\tau},A\hat{\Sigma}_{\tau}A^{\top}+Q)$ is matrix monotone for any $A,B$ of appropriate dimensions and $Q\geq 0$.
Proof.
If $\hat{\Sigma}_{\tau}\geq\Sigma_{0}$, then $\hat{\Sigma}_{\tau}-\Sigma_{0}\geq 0\Rightarrow(\hat{\Sigma}_{\tau}-\Sigma_{0})=VV^{\top}$ by the definition of positive definiteness, where V results from the Cholesky decomposition of $(\hat{\Sigma}_{\tau}-\Sigma_{0})$. Further, $A(\hat{\Sigma}_{\tau}-\Sigma_{0})A^{\top}=AVV^{\top}A^{\top}=(AV)(AV)^{\top}$ which is also, by definition, positive semi-definite. Finally $\hat{\Sigma}_{\tau+1}-\Sigma_{1}=A\hat{\Sigma}_{\tau}A^{\top}+Q-A\hat{\Sigma}_{0}A^{\top}-Q=A(\hat{\Sigma}_{\tau}-\Sigma_{0})A^{\top}\geq 0$, and thus $\hat{\Sigma}_{\tau+1}\geq\Sigma_{1}$. ∎
We can now show that the first-step scaling factor $\xi_{t+1}$ is always an upper-bound of the scaling factor required for subsequent steps $\xi_{\tau+1}$. To do this, we will compare the scaling factors resulting from performing steps 2-5 if we use $g$ as the fixed mapping converting $D_{cal}$ to $\bar{D}_{cal}$ (as done in Theorem 5.2 to obtain $\xi_{t+1}$) or if we instead use the map $h(X^{(i)})=(\hat{\mathcal{N}}(s_{t}^{(i)},\hat{\Sigma}_{\tau}),u_{t}^{(i)})$ (as must be done during planning to obtain $\xi_{\tau+1}$). Using $g$ leads to the exact scaling factor only for the first planning step, and using $h$ is computationally impractical.
Theorem 0.A.1.
If the approximate dynamics $\tilde{f}$ are linear (Assumption 0.A) and $\Sigma_{0}\leq\hat{\Sigma}_{\tau}$ (Assumption 0.A), then $\xi_{\tau+1}\leq\xi_{t+1},\forall\tau\in\{t+1,\ldots,t+H-1\}$.
Proof.
Since $D_{cal}$ is fixed a-priori, and both $g$ and $h$ (step 2) result in the same estimated state mean $\hat{\mu}_{\tau}$, then by Corollary 1, after approximate propagation (step 3) they will still share means $\hat{\mu}_{\tau+1}$. Using Assumption 0.A and Lemma 3 we further get $\hat{\Sigma}_{\tau+1}\geq\Sigma_{1}\Rightarrow\hat{\Sigma}_{\tau+1}^{-1}\leq\Sigma_{1}^{-1}$. For any $X^{(i)},Y^{(i)}$ in $D_{cal}$, let $z=s_{\tau+1}-\hat{\mu}_{\tau+1}$. We can then write $z^{\top}\hat{\Sigma}^{-1}_{\tau+1}z\leq z^{\top}\Sigma_{1}^{-1}z\Leftrightarrow d_{M}^{2}(\hat{\mathcal{N}}_{\tau+1}(\hat{\mu}_{\tau+1},\hat{\Sigma}_{\tau+1}),s_{\tau+1})\leq d_{M}^{2}(\hat{\mathcal{N}}_{\tau+1}(\hat{\mu}_{\tau+1},\Sigma_{1}),s_{\tau+1})$. Recalling the nonconformity score of Equation (12) this is equivalent to $R^{(i)}_{\tau+1}\leq R^{(i)}_{1},\forall i\in\{1,\ldots,\lvert D_{cal}\rvert\}$ where the subscript denotes the state covariance used. It follows that $\{x\in\mathbb{R}\mid R^{(i)}_{\tau+1}\leq x\}\supseteq\{x\in\mathbb{R}\mid R^{(i)}_{1}\leq x\}$ and hence $\mathbbm{1}_{\{R^{(i)}_{\tau+1}\leq x\}}\geq\mathbbm{1}_{\{R^{(i)}_{1}\leq x\}},\forall x\in\mathbb{R}$. Let us introduce $R_{\infty,\tau+1}=\{R^{(i)}_{\tau+1}\}^{\lvert D_{cal}\rvert}_{i=1}\cup\{\infty\}$ and $R_{\infty,1}=\{R^{(i)}_{1}\}^{\lvert D_{cal}\rvert}_{i=1}\cup\{\infty\}$ as in §$\S$4. By definition, $F_{R_{\infty,\tau+1}}(x)\geq F_{R_{\infty,1}}(x),\forall x\in\mathbb{R}$, where $F$ is the empirical residual distribution defined in Equation (6). From SplitCP, $\hat{q}_{\tau+1}=\mathcal{Q}(1-\alpha;F_{R_{\infty,\tau+1}})$ and $\hat{q}_{t+1}=\mathcal{Q}(1-\alpha;F_{R_{\infty,1}})$ and therefore $\hat{q}_{\tau+1}\leq\hat{q}_{t+1}\Rightarrow\frac{\hat{q}_{\tau+1}^{2}}{\mathcal{\chi}^{2}_{\dim(\mathcal{S}),\alpha}}\leq\frac{\hat{q}_{t+1}^{2}}{\chi^{2}_{\dim(\mathcal{S}),\alpha}}\Rightarrow\xi_{\tau+1}\leq\xi_{t+1}$. ∎
Hence, using $\xi_{t+1}$ as the scaling factor during planning allows us to conservatively construct calibrated regions. Finally, we show that using a larger scaling factor than required maintains the coverage guarantee.
Theorem 0.A.2.
Let $\hat{\Sigma}_{\tau+1,cal}=\xi_{t+1}\hat{\Sigma}_{\tau+1}$. If $\xi_{\tau+1}\leq\xi_{t+1}$ then $\mathbb{P}(s_{\tau+1}\in\hat{\Sigma}_{\tau+1,cal})\geq(1-\alpha)$.
Proof.
Let $\xi_{t+1}=\xi_{cons}+\xi_{\tau+1}$, where $\xi_{cons}\geq 0$. Consequently $\xi_{t+1}\hat{\Sigma}_{\tau+1}-\xi_{\tau+1}\hat{\Sigma}_{\tau+1}=\xi_{cons}\hat{\Sigma}_{\tau+1}\geq 0$. By Corollary 2, then $\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},\xi_{\tau+1}\hat{\Sigma}_{\tau+1})}\subseteq\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},\hat{\Sigma}_{\tau,cal})}$. Since by definition $\mathbb{P}(s_{\tau+1}\in\xi_{\tau+1}\hat{\Sigma}_{\tau+1})\geq(1-\alpha)$ and $s_{\tau+1}\in\xi_{\tau+1}\hat{\Sigma}_{\tau+1}\Rightarrow s_{\tau+1}\in\hat{\Sigma}_{\tau+1,cal}$ then $\mathbb{P}(s_{\tau+1}\in\hat{\Sigma}_{\tau+1,cal})\geq(1-\alpha)$.
Let us consider how the multi-step coverage guarantees derived above might hold with local calibration. Again, we can compare the results of performing steps 2-5 with fixed mappings $g$ or $h$. The partitioning is fixed prior to applying $g$ or $h$, since it was generated from $\bar{D}_{cal,part}$. Since $\mathcal{T}$ does not consider the state uncertainty, then its outputs are the same for both approaches i.e. $\mathcal{T}(g(X^{(i)}))=\mathcal{T}(h(X^{(i)})),\forall i\in\{1,\ldots,\lvert{D}_{cal,scale}\rvert\}$. Recalling the element-wise residual inequality $R^{(i)}_{\tau+1}\leq R^{(i)}_{1},\forall i\in\{1,\ldots,\lvert D_{cal}\rvert\}$, it follows that in each partition $\mathcal{X}_{k}$ we have $\hat{q}_{\tau+1,k}\leq\hat{q}_{t+1,k}\Rightarrow\xi_{\tau+1,k}\leq\xi_{t+1,k}$. Hence even with local conformal calibration $\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},\xi_{\tau+1,k}\hat{\Sigma}_{\tau+1})}\subseteq\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{\tau},\xi_{t+1,k}\hat{\Sigma}_{\tau+1})},\forall k\in\{1,\ldots,K\}$ which enables LUCCa to guarantee at least $(1-\alpha)\%$ coverage.

### v3

@@ 7.0.1 Acknowledgements
Appendix 0.A Proving LUCCa’s Multi-Step Coverage with Linear Dynamics Approximators
Our proposed algorithm (LUCCa) used MPPI [24], a sampling-based shooting-style MPC scheme, to compute “optimal” control input sequences that balance goal progression, obstacle avoidance, and uncertainty reduction - see Equation (15) and Algorithm 1. Empirically, MPPI has been shown to be effective for a wide range of approximate dynamics forms (e.g. nonlinear, non-differentiable) and constraint types (e.g. nonlinear). Consequently, the quality of MPPI’s output is highly dependent on the structure of the problem at hand, which we kept broad in the main paper, and on the available compute budget (e.g. number of sampled trajectories, planning horizon, range of control disturbances). Therefore, it is difficult to derive general claims about safety resulting from MPPI plans. To make progress, we will make the simplifying assumption that the MPPI “optimal” plan acts as a one-step feedback controller correcting for previous prediction errors. Under this assumption, if the approximate dynamics are linear-Gaussian, $\xi\geq 1$, and the system has enough control authority, we can use a procedure similar to [12] to show that the one-step errors resulting from applying the feedback law are always exchangeable with the errors observed at the first planning step. Then, by applying the conformal calibration procedure derived for the first planning step (Theorem 5.2 of the main paper), we show how we can extend the first-step safety guarantees for the remaining planning steps. We empirically observe that LUCCa can achieve higher safety rates (as shown in the experiments), since it is generally over-conservative compared to this simplified one-step feedback law.
Recall the unknown time-invariant map $f$ describing the true system dynamics, previously presented in Equation (1) of the main paper
$s_{t+1}\sim f(s_{t},u_{t})$
Without loss of generality, we will now rewrite $s_{t+1}=f(s_{t},u_{t},\eta)$, where $\eta$ is a stochastic disturbance drawn from an unknown time-invariant distribution. The approximate dynamics $\tilde{f}$ is defined to propagate a Multivariate Normal (MVN) estimate of the true state, which we denote as $\mathcal{N}(\hat{\mu}_{t+1},\hat{\Sigma}_{t+1})=\tilde{f}(\hat{\mu}_{t},\hat{\Sigma}_{t},u_{t})$. Note that both the true $f$ and approximate dynamics $\tilde{f}$ take in the same control input $u_{t}$.
Let us now define the prediction error at time $t$ as $e_{t}\coloneqq s_{t}-\hat{\mu}_{t}$, where $s_{t}$ is the true system state and $\hat{\mu}_{t}$ the expectation of the approximate dynamics prediction. Following Assumption 1 of the main paper, $e_{0}=0$ and hence we can define the first-step error as
$e_{1}$ $=s_{1}-\hat{\mu}_{1}$
$=f(s_{0},u_{0},\eta)-\mathbb{E}[\tilde{f}(\hat{\mu}_{0},\Sigma_{0},u_{0})]$
$=f(s_{0},u_{0},\eta)-\mathbb{E}[\tilde{f}(s_{0},\Sigma_{0},u_{0})]$
where $(\hat{\mu}_{0},\Sigma_{0})$ originates from the fixed mapping $g$ defined in Equation (11) and $\hat{\mu}_{0}=s_{0}$ since $e_{0}=0$. This error accounts for both model mismatch (epistemic uncertainty) and noise (aleatoric uncertainty). In the main paper, we assumed access to a dataset of first-step transitions $D_{cal}=\{(s_{0},u_{0},s_{1})^{(i)}\}_{i=1}^{\lvert D_{cal}\rvert}$ from which we computed the Mahalanobis distance between the approximate MVN $\mathcal{N}(\hat{\mu}_{1},\hat{\Sigma}_{1})$ and the true state $s_{1}$ as detailed in §$\S$5.2. Then, in Theorem 5.2, we used conformal prediction to find a scalar $\xi\in\mathbb{R}$ to calibrated the approximate MVNs, such that, averaged over the distribution of states and actions,
$\mathbb{P}(e_{1}^{\top}(\xi\hat{\Sigma}_{1})^{-1}e_{1}\leq\chi^{2}_{\dim(\mathcal{S}),\alpha})\geq(1-\alpha)$
This guarantee provided that, after the first planning step, $(1-\alpha)\%$ of the probability mass of the true states $s_{1}$ were contained in the hyperellipsoid corresponding to the $(1-\alpha)\%$ confidence region of the calibrated approximate dynamics prediction $(\hat{\mu}_{1},\xi\hat{\Sigma}_{1})=\tilde{f}_{cal}(\hat{\mu}_{0},\Sigma_{0},u_{0})$. We now aim to derive similar probability guarantees for the remainder of the planning horizon. Concretely, we aim to construct prediction regions whose $(1-\alpha)\%$ confidence ellipsoids contain $(1-\alpha)\%$ of the probability mass of the true states $s_{t}$ for $t\geq 2$. Jointly with MPC-enforced safe region constraints, this would enable us to provide probabilistic safety along the entire plan - as described by Equation (3) of the main paper.
While in the first planning step the approximate $\tilde{f}$ and true dynamics $f$ rollout from the same state $s_{0}=\hat{\mu}_{0}$, this is not necessarily the case for the following steps. In general, $s_{t}\neq\hat{\mu}_{t}$ for $t\geq 1$ and so, from the second step onwards, the approximate and true dynamics rollouts will have different starting points. This can cause an accumulation of one-step errors, resulting in $e_{t}$ for $t\geq 2$ not being necessarily exchangeable with $e_{1}$, and hence the conformal scaling factor $\xi$ being possibly insufficient to ensure coverage. If we could modify the control inputs $u_{t}$ to account for previous errors such that $e_{t}$ became exchangeable with $e_{1}$, then the first-step calibrated covariance $\hat{\Sigma}_{1,cal}$ would provide the desired coverage at every step (by Theorem 5.2). We will show how this might be achieved if $\tilde{f}$ is linear-Gaussian, the system has sufficient control authority, and the planner acts as a one-step feedback controller. {assumption} The approximate dynamics $\tilde{f}$ are linear-Gaussian, i.e.
$\mathcal{(}\hat{\mu}_{t+1},\hat{\Sigma}_{t+1})=\tilde{f}(\hat{\mu}_{t},\hat{\Sigma}_{t},u_{t})=(\tilde{A}\hat{\mu}_{t}+\tilde{B}u_{t},\tilde{A}\hat{\Sigma}_{t}\tilde{A}^{\top}+\tilde{Q})$, where $\tilde{A},\tilde{B},\tilde{Q}$ are matrices of approximate dimension and $\tilde{Q}\geq 0$. Note that the true unknown dynamics $f$ may still be nonlinear and non-Gaussian. The calibrated approximate dynamics update $\tilde{f}_{cal}$, defined in Algorithm 2, can then be rewritten as
$\hat{\mu}_{t+1}$ $=\tilde{A}\hat{\mu}_{t}+\tilde{B}u_{t}$
$\hat{\Sigma}_{t+1,cal}$ $=\xi(\tilde{A}\hat{\Sigma}_{t,cal}\tilde{A}^{\top}+\tilde{Q})$
where $\hat{\Sigma}_{0,cal}=\Sigma_{0}$ as set by the map $g$. We can now rewrite the first-step error as
$e_{1}$ $=s_{1}-\hat{\mu}_{1}$
$=f(s_{0},u_{0},\eta)-\mathbb{E}[\tilde{f}(\hat{\mu}_{0},\Sigma_{0},u_{0})]$
$=f(s_{0},u_{0},\eta)-(\tilde{A}\hat{\mu}_{0}+\tilde{B}u_{0})$
$=f(s_{0},u_{0},\eta)-(\tilde{A}s_{0}+\tilde{B}u_{0})$
Figure 5: First two planning steps with conformal calibration and the one-step feedback law. The true system evolves according to $f$ with examples of possible states shown in black. The “optimal” plan is shown in teal, along with the $(1-\alpha)\%$ confidence regions obtained from recursively applying $\tilde{f}_{cal}$. After the first planning step, if we keep applying the “optimal” control inputs $u_{t}$ in an open-loop fashion, the error (purple) between the true system state and the approximate dynamics can accumulate - case shown in light gray. Instead, by applying the “corrected” control input $\tilde{u}_{1}$ obtained via the one-step feedback law (green and red), we ensure that the error at times $t\geq 2$ remains exchangeable with the one-step error. Hence, we can maintain the original first-step coverage guarantees for the remaining planning steps when such a feedback law exists.
Similarly to [12], let us now analyze the conditions for the existence of a one-step feedback law that corrects for the drift $e_{t}$ between the planned and executed trajectories at every step. While in our “optimal” plan, each control input $u_{t}$ progresses the expectation of the approximate dynamics from $\hat{\mu}_{t}$ to $\hat{\mu}_{t+1}$, during execution, the system may be at a different state $s_{t}\neq\hat{\mu}_{t}$. A one-step feedback law would sense the current state $s_{t}$ and apply a “corrected” control input $\tilde{u}_{t}\coloneqq u_{t}+\Delta u_{t}$ that drives the planned system from $s_{t}$ to the next planned point $\hat{\mu}_{t+1}$, i.e. find $\tilde{u}_{t}$ such that $\mathbb{E}[\tilde{f}_{cal}(s_{t},\hat{\Sigma}_{t,cal},\tilde{u}_{t})]=\hat{\mu}_{t+1}\coloneqq\mathbb{E}[\tilde{f}_{cal}(\hat{\mu}_{t},\hat{\Sigma}_{t,cal},u_{t})]$. A visualization of this process is shown in Figure 5. Such a law would prevent the accumulation of error along the planned trajectory.
Since $\tilde{f}$ is linear, modifying the control input via the one-step feedback law will not change the predicted covariance $\hat{\Sigma}_{t}$. Given this objective, we can solve for the required $\Delta u_{t}$ at each step using
$\mathbb{E}[\tilde{f}_{cal}(s_{t},\hat{\Sigma}_{t,cal},\tilde{u}_{t})]$ $=\mathbb{E}[\tilde{f}_{cal}(\hat{\mu}_{t},\hat{\Sigma}_{t,cal},u_{t})]$
$\tilde{A}s_{t}+\tilde{B}\tilde{u}_{t}$ $=\tilde{A}\hat{\mu}_{t}+\tilde{B}u_{t}$
$\tilde{A}(\hat{\mu}_{t}+e_{t})+\tilde{B}(u_{t}+\Delta u_{t})$ $=\tilde{A}\hat{\mu}_{t}+\tilde{B}u_{t}$
$\tilde{A}\hat{\mu}_{t}+\tilde{B}u_{t}+\tilde{A}e_{t}+\tilde{B}\Delta u_{t}$ $=\tilde{A}\hat{\mu}_{t}+\tilde{B}u_{t}$
$\tilde{A}e_{t}+\tilde{B}\Delta u_{t}$ $=0$ (16)
If $\dim(\mathcal{U})\geq\dim(\mathcal{S})$ and $\tilde{B}$ is full rank, then we can use the Moore-Penrose pseudo-inverse of $\tilde{B}$, denoted $\tilde{B}^{\dagger}$, to obtain $\Delta u_{t}=-\tilde{B}^{\dagger}\tilde{A}e_{t}$ as the solution to the equation above. Then, we can write the one-step feedback law as
$\tilde{u}_{t}=u_{t}-\tilde{B}^{\dagger}\tilde{A}(s_{t}-\hat{\mu}_{t})$
We now make the simplifying assumption that the MPPI-computed plans correct for previous prediction errors, allowing us to make all the one-step errors exchangeable with $e_{1}$. {assumption} The “optimal” plan obtained by MPPI acts as the one-step feedback law defined above, correcting for previous prediction errors.
By applying the one-step feedback law during inference, we have that error at time $t\geq 2$ then becomes
$e_{t+1}$ $=s_{t+1}-\hat{\mu}_{t+1}$
$=s_{t+1}-\mathbb{E}[\tilde{f}_{cal}(\hat{\mu}_{t},\hat{\Sigma}_{t,cal},u_{t})]$
$=s_{t+1}-(\tilde{A}\hat{\mu}_{t}+\tilde{B}u_{t})$
$=s_{t+1}-(\tilde{A}\hat{\mu}_{t}+\tilde{B}u_{t}+\smash{\overbracket{\tilde{A}e_{t}+\tilde{B}\Delta u_{t}}^{=0}})$
$=s_{t+1}-(\tilde{A}\hat{\mu}_{t}+\tilde{A}(s_{t}-\hat{\mu}_{t})+\tilde{B}(u_{t}+\Delta u_{t}))$
$=f(s_{t},\tilde{u}_{t},\eta)-(\tilde{A}s_{t}+\tilde{B}\tilde{u}_{t})$
which has the same form as the first-step error, since the approximate and true dynamics have the same inputs. Hence, by applying the one-step feedback law, we can make $e_{t},t\geq 2$ exchangeable with $e_{1}$. From Theorem 5.2, it then follows that the $(1-\alpha)\%$ confidence region of the MVN centred at $\hat{\mu}_{t}$ with covariance $\hat{\Sigma}_{1,cal}=\xi\hat{\Sigma}_{1}$ will contain at least $(1-\alpha)\%$ of the probability mass of the true states $s_{t}$. This allows us to extend LUCCa’s one-step safety guarantees to be valid across multiple steps.
Theorem 0.A.1.
Let $\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{t},\hat{\Sigma}_{1,cal}),\alpha}$ denote the $(1-\alpha)\%$ confidence hyperellipsoid of the MVN centered at $\hat{\mu}_{t}$ with covariance $\hat{\Sigma}_{1,cal}$. If the planner acts as a one-step feedback controller, $\dim(\mathcal{U})\geq\dim(\mathcal{S})$, and $\tilde{B}$ is full rank, then the true states $s_{t}$ are contained in the $(1-\alpha)\%$ confidence region of $\mathcal{N}(\hat{\mu}_{t},\hat{\Sigma}_{1,cal})$ with probability at least $(1-\alpha)$ at each time-step $t\in\{1,\ldots,T\}$.
@@ Appendix 0.A Proving LUCCa’s Multi-Step Coverage with Linear Dynamics Approximators
For $t=1$, this is ensured by conformal prediction as shown in Theorem 5.2 of the main paper. For $t\geq 1$, if $\dim(\mathcal{U})\geq\dim(\mathcal{S})$ and $\tilde{B}$ is full rank then $\tilde{u}_{t}$ exists and is exact. By sensing $s_{t}$, we can calculate $\tilde{u}_{t}$ at every step. If the planner acts as a one-step feedback controller producing $\tilde{u}_{t}$, we can apply the input to the true and approximate systems, correcting for $e_{t}$ and making $e_{t+1}$ exchangeable with $e_{1}$. Then by Theorem 5.2, it follows that $\mathbb{P}(s_{t+1}\in\varepsilon_{\hat{\mathcal{N}}(\hat{\mu}_{t+1},\hat{\Sigma}_{1,cal}),\alpha})\geq(1-\alpha)$. ∎
Recall that LUCCa’s calibrated predicted covariance at time $t$ is $\tilde{\Sigma}_{t,cal}=\sum_{i=0}^{t}\xi_{1}^{i}\tilde{A}^{i}\Sigma_{0}(\tilde{A}^{i})^{\top}+\sum_{i=0}^{t-1}\xi_{1}^{i+1}\tilde{A}^{i}\tilde{Q}(\tilde{A}^{i})^{\top}$, which is generally overly conservative relative to the one-step calibrated covariance $\hat{\Sigma}_{1,cal}$ if $\xi\geq 1$. This conservative compounding of calibration factors was chosen to mitigate the impact of the planner possibly not acting as an exact one-step feedback controller. If the system has control input bounds, then one might also need to consider whether $\tilde{u}_{t}$ is admissible at each point along the plan. The proof above used a single global scaling factor $\xi$. We leave an extension guaranteeing multi-step local coverage for future work.